if the value of base is (x+16) inches and the height is (x+9) inches. the area of the triangle is 130 inches. then what is the value of x in the base and in the height?

Answer by mananth(16946)

Area of triangle = 1/2 * base*height
Area = 1/2 * (x+16)(x+9)
130= 1/2 * x^2+25x+144
multiply by 2
260= x^2+25x+144
x^2+25x-116=0
x^2+29x-4x-116=0
x(x+29)_4(x+29)=0
(x-4)(x+29)=0
x= 4 OR -29.
Ignore negative value.
x= 4